solve this the current i in a circuit is inversely proportional to its resistance R measured in ohms. suppose that when the current in a circuit is 30 amperes the resistance is 8 ohms. find the current in the same circuit when the resistance is 10 ohms
Answer by Alan3354(69443)

I = V/R (V is the constant in this, the Voltage)
30 = V/8
V = 240
----
I = V/R = 240/10
I = 24 amps